Solve by the elimination method.
4x-9y=6.5
7y-2x=-4.5

Answer by Alan3354(69443)

4x-9y=6.5
7y-2x=-4.5 times 2 --> -4x+14y = 9
-------------
+4x -9y = 6.5
-4x+14y = 9
-------------- Add
5y = 15.5
y = 3.1
---------
4x - 9y = 6.5
4x - 9*3.1 = 6.5
4x = 34.4
x = 8.6